OpenStudy (anonymous):

The function f(x) is shown below: f(x) = (3)^x - 2 The function f(x) is shifted to the left by 13 units. Which of the following best represents the new function? f(x) = (3)x + 11 f(x) = (3)x + 13 f(x) = (3)x -15 f(x) = (3)x - 13

OpenStudy (anonymous):

f(x) = (3)^x + 11 f(x) = (3)^x + 13 f(x) = (3)^x -15 f(x) = (3)^x - 13

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

Someone please help!

OpenStudy (anonymous):

here are the rules for shifting horizontally f(x) f(x+c) shifts it left. SO...

OpenStudy (anonymous):

oh okay, well when I did it, I got f(x) = (3)x + 11 ?

OpenStudy (anonymous):

yeah, but wouldn't it change to 11 since the 2 is already in the equation?

OpenStudy (anonymous):

wait you wrote it weird use perenthesis

OpenStudy (anonymous):

well all of the answers are cubed

OpenStudy (anonymous):

(3)^x - 2 in my mind and to a computers mind it is read as 

OpenStudy (anonymous):